ou arrange two identical charges along the x-axis at positions (-2.0,0) and (2.0,0), with distances measured in centimeters. You've set things up so that the origin of your coordinates lies at the midpoint between the charges. You measure the electric force acting on the charge on the right and find it has a magnitude of 83082.0 N. You want to deduce the electric field vector that is acting on the right-hand charge. [Note that there are three answer boxes: enter the x-component of the field in the first box, the y-component of the field in the second box, and the z-component in the third box.]
Answer #1

Let the charge be q @ x = -2.0 cm = - 2.0 * 10^-2 m
Let the charge be q @ x = 2.0 cm = 2.0 * 10^-2 m

Electric force on the right charge = K * q *q / r^2
83082.0 = (8.9 * 10^9)/ (4*10^-2)^2 * q^2
q = sqrt((83082.0 * (4*10^-2)^2) / (8.9 * 10^9))
q = 1.22 * 10^-4 C

We know,
F = q * E
83082.0 = 1.22 * 10^-4 * E
E = 83082.0 * 10^4 / 1.22 N/C
E = 6.81 * 10^8 i N/C

X component = 6.81 * 10^8
Y component = 0
Z component = 0